Using polar coordinates, evaluate the integral ∫∫Rsin(x^2+y^2)dA where R is the region 16≤x^2+y^2≤25.
1
Expert's answer
2020-11-02T18:38:52-0500

"\iint_R \sin(x^2+y^2)dA, R=\{16\leq x^2+y^2\leq 25\}"


"x = r\cos \phi \\y = r \sin\phi\\ I = r\\"

Find the integration limits:

"\phi \in [0,2\pi]"

"16 \leq x^2 + y^2 \leq 25\\
16 \leq r^2\cos^2\phi + r^2\sin^2\phi \leq 25\\
16\leq r^2\leq 25\\
4\leq r \leq 5"

"r \in[4,5]"

So:

"\iint_R sin(x^2 + y^2)dA = \int_4^5dr\int_0^{2\pi} rsin(r^2)d\phi = \\
=\int_4^5rsin(r^2)dr \phi|_0^{2\pi} = 2\pi\int_4^5 rsin(r^2)dr = \\
=
\left[
 \begin{array}{ccc}
 r^2 = u & 5 \to 25\\
 2rdr = du& 4 \to 16\\
 \end{array}
\right] = \pi \int_{16}^{25}sin(u)du =\\
=-\pi cos(u)|_{16}^{25} =\pi(cos16 - cos25) = \\=2\pi sin(\frac{41}{2})sin(\frac{9}{2})"